Ask someone borrowing from a five-six ( Bumbay) money lender. How much w be charged if you want to loan P5,000.00, payable in one year? Determine the interest rate per period and the annual interest rate. Is it a good loan term? Whi or why not?




Submit on a long bond paper an essay containing your answer on the questions above. You should note that the computation must be included in the front pag of your paper. Then essay must consist of at least 5 sentences and a maximum of 10 sentences on why you must chose your answer. Justify your answer with the computations you had and your own personal reasons.

1
Expert's answer
2022-02-07T15:13:55-0500

SOLUTION:

=5000*10/100*1

= Rs 500 is the interest.

Explanation:

Interest is the money you pay to money lender when you borrow money be it a bank. If you borrow a Rs 5000 personal loan, you may wind up paying the lender a total of over Rs 5000 in the next one year. That extra cost is the interest.

If a lender uses the simple interest method , it is easy to calculate the interest on he loan.

If the loan taken was for Rs5000 and the interest rate on the loan is 10 %, the simple interest formula is as follows:

5000*10/100*1 = 500 as the interest

But in real life money lender calculates as compound interest. If you are paying annually for one year, it will be the same as simple interest. If you are paying for more than year it will change hence to be calculated using the formula:

A=P(1+r/100)*n